OpenStudy (anonymous):

Calculate the number of moles in 0.457 grams of NaCl, sodium chloride.

OpenStudy (anonymous):

0.0078

OpenStudy (anonymous):

Explain why the number you calculated for Avogadro’s number may not match theactual number of 6.02 x 10^23

OpenStudy (ash2326):

atomic mass of Na Cl= 23.5+35=58.5 amu 1 mole weighs 58.5 gm so 1 gram has 1/58.5 moles 0.457gm will have 0.457/58.5=0.0078 moles

OpenStudy (ash2326):

1 mole contains 6.023*10^23 molecules here we have 0.0078 moles so 0.0078*6.023*10^23= 4.6974*10^6 mlecules

OpenStudy (anonymous):

\[n=M/M _{a}\] n=0.457/(23+35)=7.9*10^(-3)

OpenStudy (ash2326):

Hayat Cakes, next time post chemistry questions in chemistry group. Moderator can delete your post.

OpenStudy (anonymous):

oh okay thanks.

OpenStudy (ash2326):

anyways, did you get the problem